# TENSLORA: TENSOR ALTERNATIVES FOR LOW-RANK ADAPTATION

### v1

@@ 1 Introduction
Despite its strengths, LoRA applies updates independently across attention projections (Query, Key, and Value) and layers, potentially disregarding strong correlations along these dimensions and thus creating undesired redundancy, and scaling linearly with the number of adapted components. Several tensor-based extensions, such as FacT [5], LoTR [6], LoRTA [7], and CaRA [8], attempt to reduce redundancy by aggregating LoRA parameters into higher-order tensors. However, these approaches explore only partially ways to extend LoRA to tensors and provide little guidance on which choices are the most effective.
In this work, we introduce TensLoRA, a unified framework that systematically covers tensor-based low-rank adaptations by aggregating attention layers into tensors. TensLoRA captures and extends FacT, LoTR, LoRTA, and CaRA, albeit with differences in scope and implementation, particularly with respect to the types of layers considered. Beyond unification, TensLoRA introduces mode-specific compression rates, offering finer control over parameter budgets and enabling compression rates tailored to the characteristics of specific data or tasks. Our experiments show that performance depends on the tensor construction, with some variants surpassing standard LoRA under comparable parameter counts, though none consistently outperforms LoRA under heavy compression rates. Code to reproduce experiments can be found on GitHub  [^1] https://github.com/ax-le/TensLoRA.
@@ 2.2.1 Intra-Attention modules — Att
Transformers use multi-head attention, where each attention layer is composed of $h$ heads. Each head typically operates on a projection of size $d_{h}=d/h$, i.e. a subset of the inner dimension. In most LoRA implementations, a single low-rank adaptation is shared across all heads. In the Att variant, we propose to separate the LoRA adaptation across heads, based on the hypothesis that different heads may be adapted using similar subspaces. This hypothesis may be too strong, because some previous work has suggested that each attention head holds a specialized role [9].
@@ 2.2.3 Depth of the model — Depth
Stacking the same type of projection (e.g., Query matrices) from each of the $L$ layers results in a tensor $\mathcal{W}_{0}~\in~\mathbb{R}^{d\times d\times L}$, as presented in Figure 1(c). Depth is based on the hypothesis that different layers in the model may be adapted using similar subspaces.
@@ 3 Tensor Factorization
where $\mathcal{G}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ is called the core tensor, $A\in\mathbb{R}^{a\times r_{1}},B\in\mathbb{R}^{b\times r_{2}},C\in\mathbb{R}^{c\times r_{3}}$ are the factor matrices along the first, second, and third modes, respectively. The operator $\times_{n}$ denotes the mode-$n$ tensor-matrix product. The number of factor matrices is equal to the number of modes of the original tensor, and the core tensor has the same number of modes as the original tensor.
@@ 4.1 Datasets
We investigated two datasets for both modalities (Vision and Language): DTD [17] and EuroSAT [18] for Vision, and CoLA and MRPC for Language, both part of the GLUE benchmark [19]. These datasets were used in related work (FacT [5] and CaRA [8] for Vision, and LoTR [6] and LoRTA [7] for Language). All these datasets are used in a classification setting.
@@ 4.2 Experimental and Model details
During our experiments, we noticed that two parameters were largely impacting results: the scaling factor $\alpha$ (as presented in [6], also used in our implementation of LoRA) and the initialization of tensors. In this paper, based on preliminary experiments, we fixed $\alpha=4$, and the initialization of tensors to “orthogonal,” thus resembling HOSVD [16].
@@ 4.3 Rank Selection
We consider two strategies for selecting tensor ranks, representing two limit cases relative to LoRA: “isorank” and “isoparameters”. In isorank, representing the lowest parameter count, all modes share the same rank, set equal to the LoRA rank. In isoparameters, ranks are chosen so that TensLoRA matches the number of parameters in LoRA, hence resulting in the upper bound. Since multiple rank configurations could satisfy this constraint, we adopt a simple heuristic: the QKV dimension is fixed to rank 3 and Att and Depth to 12, corresponding exactly to the sizes of these modes. The two remaining ranks are then distributed such that modes of equal size share the same rank, and larger modes receive larger ranks, while keeping the total parameter count close to LoRA.
@@ 4.4 Experimental results
Experimental results are presented in Table 2 where LoRA rank is fixed to 4. We present results as the average and the standard deviation over five runs.
QKV_Depth and Att_QKV_Depth consistently outperform LoRA under the isoparameters condition, establishing them as our most promising methods. This result validates the design choices seen in related works, as they employ similar tensor constructions. It also quantitatively confirms that leveraging tensors may improve LoRA, given the same number of parameters. Interestingly, these are the only two conditions that combine all LoRA matrices into a single tensor rather than several, independent ones.
@@
In particular, the Att method proves to be the least effective, almost always yielding the worst results in our isoparameters tests. In isorank tests, the QKV method outperforms Att on both EuroSAT and MRPC datasets despite having fewer parameters. A similar trend is observed where QKV_Depth outperforms Att_Depth. These findings suggest that the QKV and depth dimensions are the most critical to group in tensors, likely due to higher redundancy.
@@ 5 Conclusion
In this work, we introduced TensLoRA, a unified framework that generalizes low-rank adaptation into tensor-based formulations. By systematically aggregating LoRA updates into higher-order tensors and applying Tucker factorization, we highlighted a wide spectrum of possible adaptations that extend and encompass prior approaches. Our experiments show that adaptations do not perform equivalently, suggesting that redundancy across modes is not uniform. In particular, aggregating over the attention heads dimension appears less efficient than the other dimensions (attention projection and depth of the model). Although LoRA remains difficult to outperform with drastically fewer parameters, our analysis highlights promising tensor structures for future exploration.
Future research should further investigate rank allocation strategies, notably ablation studies over different dimensions, explore alternative factorizations (CP, Tensor-Train), and extend the framework to other modules, such as output projections or the MLP layers. Additionally, we suggest that TensLoRA may serve as a tool for interpretability, enabling the analysis of redundancies across modes and the identification of critical components in Transformers.

### v2

@@ 1 Introduction
Despite its strengths, LoRA applies updates independently across attention projections (Query, Key, and Value) and layers, potentially disregarding correlations along these dimensions and creating undesired redundancy; furthermore, the number of parameters scales linearly with the number of adapted components. Several tensor-based extensions, such as FacT [13], LoTR [2], LoRTA [11], and CaRA [21], attempt to reduce redundancy by aggregating LoRA parameters into higher-order tensors. However, these approaches explore only partially ways to extend LoRA to tensors and provide little guidance on which choices are the most effective.
In this work, we introduce TensLoRA, a unified framework that systematically covers tensor-based low-rank adaptations by aggregating attention layers into tensors. TensLoRA captures and extends FacT, LoTR, LoRTA, and CaRA, albeit with differences in scope and implementation, particularly with respect to the types of layers considered. Beyond unification, TensLoRA introduces mode-specific compression rates, allowing parameter budgets to be strategically allocated to exploit non-uniform redundancy across different model dimensions. Our experiments show that performance depends on the tensor construction, with some variants surpassing standard LoRA at comparable parameter counts, though none consistently outperforms LoRA at high compression rates. Code to reproduce experiments can be found on GitHub  [^1] https://github.com/ax-le/TensLoRA/tree/ICASSP26.
@@ 2.2.1 Intra-Attention modules — Att
Transformers use multi-head attention, where each attention layer is composed of $h$ heads. Each head typically operates on a projection of size $d_{h}=d/h$, i.e. a subset of the inner dimension. In most LoRA implementations, a single low-rank adaptation is shared across all heads. In the Att variant, we propose to separate the LoRA adaptation across heads, based on the hypothesis that different heads may be adapted using similar adaptation subspaces. This hypothesis may be too strong, because some previous work has suggested that each attention head holds a specialized role [8].
@@ 2.2.3 Depth of the model — Depth
Stacking the same type of projection (e.g., Query matrices) from each of the $L$ layers results in a tensor $\mathcal{W}_{0}~\in~\mathbb{R}^{d\times d\times L}$, as presented in Figure 1(c). Depth is based on the hypothesis that different layers may be adapted using similar subspaces.
@@ 3 Tensor Factorization
where $\mathcal{G}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ is called the core tensor, $A~\in~\mathbb{R}^{a\times r_{1}}$, $B~\in~\mathbb{R}^{b\times r_{2}},C~\in~\mathbb{R}^{c\times r_{3}}$ are the factor matrices along the first, second, and third modes, respectively. The operator $\times_{n}$ denotes the mode-$n$ tensor-matrix product. The number of factor matrices is equal to the number of modes of the original tensor, and the core tensor has the same number of modes as the original tensor.
@@ 4.1 Datasets
We evaluated our methods on four datasets across two modalities (Vision and Language): DTD [5] and EuroSAT [10] for Vision, and CoLA and MRPC for Language, both part of the GLUE benchmark [22]. These datasets were used in related work (FacT [13] and CaRA [21] for Vision, and LoTR [2] and LoRTA [11] for Language). All these datasets are used in a classification setting.
@@ 4.2 Experimental and Model details
During our experiments, we noticed that two parameters were largely impacting results: the scaling factor $\alpha$ (as presented in [2], also used in our implementation of LoRA) and the initialization of tensors. In this paper, based on preliminary experiments, we fixed $\alpha=4$ and the initialization of tensors to “orthogonal,” thereby resembling HOSVD [6].
@@ 4.3 Rank Selection
We consider two strategies for selecting the tensor ranks $r_{i}$, representing two limit cases relative to LoRA: “isorank” and “isoparameters”. In isorank, representing the lowest parameter count, all modes share the same rank, set equal to the LoRA rank. In isoparameters, ranks are chosen so that TensLoRA matches the number of parameters in LoRA, yielding an upper bound. Since multiple rank configurations could satisfy this constraint, we adopt a simple heuristic: the QKV dimension is fixed to rank 3 and Att and Depth to 12, corresponding exactly to the sizes of these modes. The two remaining ranks are then distributed so that modes of equal size share the same rank, and larger modes receive higher ranks, while keeping the total parameter count close to that of LoRA. We employ this simple heuristic to avoid overfitting ranks, aiming to isolate the impact of the tensor structure.
@@ 4.4 Experimental results
Experimental results are presented in Table 2 where the LoRA rank $r$ is fixed to 4. We present results as the average and the standard deviation over five runs.
QKV_Depth and Att_QKV_Depth consistently outperform LoRA under the isoparameters condition, establishing them as our most promising methods. This result validates the design choices made in related works, which employ similar tensor constructions. It also quantitatively confirms that leveraging tensors can improve LoRA performance with the same number of parameters. Interestingly, these are the only two conditions that combine all LoRA matrices into a single tensor rather than several, independent ones.
@@
In particular, the Att method proves to be the least effective, almost always yielding the worst results in our isoparameters tests. In isorank tests, the QKV method outperforms Att on both EuroSAT and MRPC datasets despite having fewer parameters. A similar trend is observed where QKV_Depth outperforms Att_Depth. These findings suggest that the QKV and depth dimensions are the most critical to group in tensors. This is likely due to higher redundancy, hence indicating that projections and layers share similar adaptation subspaces, whereas attention heads maintain specialized roles that degrade when forced into a shared representation.
@@ 5 Conclusion
In this work, we introduced TensLoRA, a unified framework that generalizes low-rank adaptation into tensor-based formulations. By systematically aggregating LoRA updates into higher-order tensors and applying Tucker factorization, we highlighted a wide spectrum of possible adaptations that extend and encompass prior approaches. Our experiments show that adaptations do not perform equivalently, suggesting that redundancy across modes is not uniform. In particular, aggregating over the attention heads dimension appears less efficient than the other dimensions (attention projection and depth of the model). Although LoRA remains difficult to outperform with drastically fewer parameters, our analysis highlights promising tensor structures for future exploration, with some outperforming LoRA in similar parameter budgets.
Future research should investigate more thorough rank allocation strategies, notably via ablation studies across different dimensions, explore alternative factorizations (CP, Tensor-Train), and extend the framework to other modules, such as output projections and MLP layers. Additionally, we suggest that TensLoRA may serve as a tool for interpretability, enabling the analysis of redundancies across modes and the identification of critical components in Transformers.